40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed May 22 15:07:09 2019, awade, DailyProgress, WOPO, Demodulation and subtraction for homodyne detector with Zurich box IMG_6099.JPGSweep_DemodLPFValues.pdfSingleHDChannel_RelaxOscillations_CompairNoiseEaterONOFF.pdfDemodFreqSweep_median100Hzto1kHz_vsdemodFreq.pdfCompairSubtractedSignalToDarkNoise.pdf
    Reply  Thu May 23 15:29:58 2019, awade, DailyProgress, WOPO, Making time domain measurement of noise power in homoydne with WOPO pump on RMSFunctionTimeCompairingSNWithInjected532nmpump.pdf
       Reply  Sun May 26 21:24:38 2019, awade, DailyProgress, WOPO, Excess noise when pumping: checklist for Monday. 
          Reply  Wed May 29 15:24:33 2019, awade, DailyProgress, WOPO, Update to checklist 20190528_ProperIQReconstructionAndLPFilteringData.zip
Message ID: 2356     Entry time: Wed May 29 15:24:33 2019     In reply to: 2354
Author: awade 
Type: DailyProgress 
Category: WOPO 
Subject: Update to checklist 

I had a look at a few of these things.  I've found that it doesn't seem to be caused by heating of the beam splitter, while turning 532 nm on an off I see no change in the balancing of the beam splitter.  It appears that on remeasuring the data on Tuesday I found that there had been some glitching of the digitized data readout.  

Suspicions for now is that there is not as much pumping light making it to the chip as a thought and that bandwidth resolution of the time scan scan was a little too wide.  I have shortened the scan, increased the noise bandwidth of the demodulation and will widen the width of the low pass filter on the noise-time scan to increase the time resolution.

I'm also implementing proper subtraction of the signal that uses both the I and Q quadratures so that all the information about the relative phase and amplitude of the digitized signals is properly mixed down by the FPGA inside the Zurich.

Notebook with demo of the subtraction is attached in a zip below.

Quote:

There are a few other possibilities for the excess noise when injecting pump, this is a checklist for me to run through tomorrow:

  1. Its possible that the waste pump light from the 532 nm is heating the fiber beam splitter a bit.  This might cause imbalance when I go to measure the noise level for pumped WOPO.  I checked both channels detector on the oscilloscope with 532 nm turned on and off.  I couldn't discern a change in the balancing​.  I will need to look closer tomorrow with the actual zurich demodulated data.  Task for tomorrow is to compare the BS balancing with 532 nm on and off.  This can be done by tuning the WOPO temperature down by 15 C (to put it outside the phase matching region) and running the SN measurement with 532 nm on and off;  Doesn't appear to be true I don't see any change in ballencing once I looks more carefully with 532 nm pumping on and off.
  2. The excess wasted 532 nm my also be locally heating the SPDC chip, so the ideal phase matching temperature might not be right.  I will try a series of temperature steps much smaller than the FWHM of the phase matching sinc curve and see if turning the temperature down/up makes a difference; I stepped through temperatures from the ideal set point of 60.99 C down to 58.00 C I saw some possible anti-squeezing peaking at 60.30C but was unable to reproduce this result.  I'll try again once I've implemented the quadrature subtraction in python dict form.
  3. Check the scanning of the 532 nm PZT is actually working.  Loop the patch cable back to the launch of the light into the fiber and use some of the dumped light from the 532 nm power control to make an interference measurement.  Here a large fringe visibility is not super important.  Just need to be able to count fringes/volt;
  4. Need to verify how much light is making it into the WOPO.  This is difficult to do directly.  But a quick check is to see how much is making it out the other side.  The PM 1064 nm fiber should carry the 532 nm pump light out the other side of the WOPO without too much loss.  Check with power meter how much 532 nm is coming out other end of WOPO;  I see 21 mW at the output of the patch cable.  There is visibly a lot of waste light exiting the fiber through the cladding.  Its not clear how much is being lost.  It looks bright to the eyes but that doesn't mean its all getting lost there.
  5. Need to document the level of power fluctuations of the 1064 nm LO light in the homodyne.  This could be leading to the fluctuating power level that is present after upping the demodulation band width and boxcar window.  If this is an issue I need to figure out quickly how to up the noise band width so that the boxcar window can be made very small and data capture is faster that the power variations of the 1064 nm.  If we had some DC monitor we might be able to correct this out in post processing;
  6. I could just be smearing anti squeezing over both measurement quadratures.  Need to check my measurement sweep times and confirm my boxcar window for scanning noise as a function of time is narrow enough;  I'm fixing this with the update to quadrature subtraction, widening demodulation noise bandwidth and implementing proper LP scipy.signal.filtfilt filtering instead of the bad box car method.
  7. Consider if adding some minicircuits pre-amps before zurich would make measurment of SN -> SQZ clearance clearer. Adding pre-amps is still a good idea, just not absolutely necessary at this point

 

Attachment 1: 20190528_ProperIQReconstructionAndLPFilteringData.zip  25.793 MB  Uploaded Thu May 30 15:04:26 2019
ELOG V3.1.3-